OpenStudy (anonymous):

√x + 2 = 2 then (x + 2)² =

OpenStudy (anonymous):

see this questions is following: first square both sides we get

OpenStudy (anonymous):

\[(\sqrt{x}+2)^{2}p =2^{2}\]

OpenStudy (anonymous):

p = 2sorry

OpenStudy (anonymous):

ohk now we know (a+b)^2 = \[a^{2}+2ab +b ^{2}\]

OpenStudy (anonymous):

\[hence, (\sqrt{x}^{2}+2(\sqrt{x}+2)+2^{2} = 4\]

OpenStudy (anonymous):

now we get \[x +2\sqrt{x}+4+4=4\]

OpenStudy (anonymous):

\[x +2\sqrt{x} = 4\]

OpenStudy (anonymous):

now we will solve (x=2)^2

OpenStudy (anonymous):

\[x ^{2}+4x +4\]

OpenStudy (anonymous):

i will continue my answer after some time i have to go seriously sorry

OpenStudy (anonymous):

Thank you for all your help,